3. A monopolist is able to practice third-degree price discrimination between two markets. The demand function in the first market is q = 500 - 2p and the demand function in the second market is q = 1,500 - 6p. To maximize his profits assuming constant marginal cost, he should

a. charge a higher price in the second market than in the first. b. charge a higher price in the first market than in the second. c. charge the same price in both markets.
d. sell only in one of the two markets.

e. None of the above.
Answer #1

b. charge a higher price in the first market than in the second.

Explanation:
q = 500 - 2p
Horizontal intercept: When p = 0; q = 500
Vertical intercept: When q = 0; 500 - 2p = 0
So, 2p = 500
So, p = 500/2 = 250

q = 1,500 - 6p
Horizontal intercept: When p = 0; q = 1500
Vertical intercept: When q = 0; 1500 - 6p = 0
So, 6p = 1500
So, p = 1500/6 = 250

So, their vertical intercept is same and horizontal intercept in first market is lesser than the horizontal intercept in second market. So, demand curve in first market is steeper which means it is less elastic and demand in second market is more elastic. To maximize profits, higher price should be charged in market with less elastic demand (market 1) and lower price in market with more elastic demand (market 2).